if a= 2x^2i-3yzj+xz^2k then div a is ?


1
Expert's answer
2021-08-27T14:16:41-0400

"a = 2x^2i-3yzj+xz^2k
\\diva= \nabla . F
\\= \frac{\partial}{\partial x}2x^2 - \frac{\partial}{\partial y} 3yz + \frac{\partial}{\partial z}xz^2
\\4x-3z+2xz"